OpenStudy (anonymous):

Divide. Write each quotient in simplest form. 2/3 ÷ 2 = _____ can somone please teach me how to do this!??

OpenStudy (anonymous):

\[\frac{ 2 }{ 3 } \div \frac{ 2 }{ 1 } = \frac{ 2 }{ 3 } \times \frac{ 1 }{ 2 } =\frac{ 2 }{ 6 }=\frac{ 1 }{ 3 }\]